# Mesh-Informed Neural Operator : A Transformer Generative Approach

### v2

@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Science Foundations for Energy Earthshot under Award Number DE-SC0024705. ZER is supported by a fellowship from the David and Lucile Packard Foundation.
@@ A.1 Datasets description
Shallow Water. This dataset contains solutions to the shallow-water equations for a 2D radial dam-break scenario on a square domain, from PDEBench (Takamoto et al., 2024). Each of the 1,000 simulations has 1,000 time steps at $128\times 128$ resolution; we downsample spatially to $64\times 64$ for efficiency and treat each time step as an independent snapshot. We randomly select 30,000 snapshots for training and 5,000 for testing
Darcy Flow. This dataset contains steady-state solutions of 2D Darcy Flow over the unit square, obtained directly from the PDEBench benchmark (Takamoto et al., 2024). We downsample the original $128\times 128$ resolution to $64\times 64$. The dataset contains 10,000 samples and we split it into 8,000 samples for training and 2,000 for testing.
@@
Global Climate. We use the real-world global climate dataset from (Dupont et al., 2022), which contains global temperature measurements over the last 40 years. Each data sample is a function defined on a grid of $46\times 90$ evenly spaced latitudes and longitudes. Follow the previous pipeline (Dupont et al., 2022), we convert the latitude-longitude pairs to Euclidean coordinates ($\mathbb{R}^{3}$) before passing them to the models. The dataset contains 9,676 training samples and 2,420 test samples.
@@ A.2 Baselines description
For all baseline models, we adopt their official implementations to ensure reproducibility; the corresponding code repositories are linked in the referenced papers. The key initialization hyperparameters for the baselines : GINO (Li et al., 2023), UPT (Alkin et al., 2025), Transolver (Wu et al., 2024), LNO (Wang & Wang, 2024), and FNO (Li et al., 2021)—are detailed in Table 6. Note that a slightly different configuration was used for the Global Climate experiment. For a complete understanding of all arguments, we refer the reader to the official repository for each respective model (github link provided in all referred papers)
@@
Latent Grid $32\times 32$
@@ A.3 Hyperparameters for MINO
MINO-T. The GNO maps input functions to a latent representation on a $16\times 16$ grid of query points, defined over the $[0,1]^{2}$ domain. We set the GNO search radius to $0.07$, the latent dimension $L_{\text{dim}}$ to 256, and the number of attention heads to 4. The encoder consists of $M_{1}=5$ cross-attention blocks and the decoder has $M_{2}=2$ blocks. For the Global Climate dataset, which is defined on a spherical manifold $\mathbb{S}^{2}$, we adjust the latent query positions to a $32\times 16$ spherical grid and increase the GNO radius to $0.2$ to account for the different coordinate system; other hyperparameters remain unchanged. The total parameter count for MINO-T is 21.5 M.
MINO-U. The GNO maps input functions to a latent representation on a $16\times 16$ grid of query points, defined over the $[0,1]^{2}$ domain. We set the GNO search radius to $0.07$, the latent dimension $L_{\text{dim}}$ to 256, and the number of attention heads to 4. The encoder consists of $M_{1}=2$ cross-attention blocks and the decoder has $M_{2}=2$ blocks. For its latent-space processor, we adopt a Diffusion U-Net architecture from the torchcfm library (Tong et al., 2024), which operates on the $[16,16]$ latent tensor with 64 channels, 1 residual block, and 4 attention heads for the processor. When applied to the Global Climate dataset, the number of U-Net channels is increased to 96 to accommodate the larger $32\times 16$ latent grid, while other settings remain the same. This configuration results in a total model size of 26.9 M for Global Climate dataset and 19.2 M for others.
@@ A.4 Details for training and inference
Reference Measures. We choose $\mu_{0}$ as a Gaussian measure characterized by a Gaussian Process (GP) with a Matérn kernel. Unless otherwise specified, the function domain is $[0,1]^{2}$ (or subset of it), and we use a kernel length scale of $0.01$ and a smoothness parameter of $0.5$. For the Global Climate dataset, the domain is a spherical manifold represented as a subset of $[-1,1]^{3}$, and the kernel length scale is adjusted to $0.05$. Furthermore, for the GP on the sphere, we explored two distance metrics for the kernel. The first uses the chordal distance (the Euclidean distance in the $\mathbb{R}^{3}$ embedding space), while the second employs the geodesic distance on $\mathbb{S}^{2}$, for which we leverage the GeometricKernels library (Mostowsky et al., 2024). Although both approaches yield a valid GP, the results presented in this paper are based on the chordal distance implementation. We selected this method for two primary reasons: its superior computational speed and to ensure consistency with other benchmarks. To facilitate further exploration, we have made both implementations publicly available in our code repository. The variance is fixed at 1 for all GPs.
Training Details. We train all models for 300 epochs using the AdamW optimizer with an initial learning rate of 1e-4. We employ a step learning rate scheduler that decays the learning rate by a gamma of $0.8$ every 25 epochs. The default batch size is 96. However, Transolver consumes significantly more GPU memory than other models. To accommodate it on a single NVIDIA RTX A6000 Ada GPU (48 GB memory), we reduced its batch size to 48 and, to maintain a comparable training iteration (duration), limited its training to 200 epochs. Despite these adjustments, Transolver still required approximately 1.76x and 1.53x more total GPU-hours than MINO-T and MINO-U, respectively. The only exception was the Global Climate experiment, where we explicitly matched the total GPU computation time of MINO variants to that of Transolver by extending their training epochs to make a fair comparison. All experiments reported in Table 4 and Table 3 were conducted three times, and we report the best performance (among the three) achieved for each model.
Inference Details. To evaluate the models, we generate the same number of samples as contained in the test set for each dataset shown in Table 2. All samples are generated by solving the learned ODE numerically using the dopri5 solver from the torchdiffeq library (Chen et al., 2019), with an error tolerance set to 1e-5 for all experiments.
@@ Appendix B SWD and MMD as general metrics for functional generation tasks
Sliced Wasserstein Distance (SWD). In practice, we estimate the distance between two probability measure,$\mu$ and $\nu$, using the discretized function samples in the datasets $X$ and $Y$. To make the metric shown in Eq. 9 computable, we need to derive the discretized version of it, which is achieved by choosing the base measure $\gamma_{s}$ to be the Haar measure (uniform surface measure) on the sphere $\mathbb{R}^{d-1}$ ($d$ is the discreziation of functions and $d=N_{in}$ if co-domain is $1$ in our case) and replace the integral and normalization with an expectation with respect to the probability measure defined by normalizing $\gamma_{s}$. Finally, we can get a discretized version of Eq. 9:
@@
where $\theta$ is a random direction on the unit sphere $\mathbb{S}^{d-1}$. SWD converges to the exact p-Wasserstein distance for two probability measures as the number of projections goes to infinity. In our experiments, we use the official implementation from the POT library (Flamary et al., 2021). We use the Sliced Wasserstein-2 distance ($p=2$(Bonneel et al., 2015), the time complexity for SWD is $\mathcal{O}(L\cdot N\cdot N_{in}+L\cdot N\log N)$, where $L$ is the number of projections.
@@
To validate this approach, we created two distinct Gaussian measures, characterized by GP, $\text{GP}_{1}$ (length scale=0.3, smoothness=1.5) and $\text{GP}_{2}$ (length scale=0.01, smoothness=0.5), on a $64\times 64$ regular grid. We drawn 5,000 GP samples from each measure to create $X,Y$ and then calculated the averaged-SWD between them for varying $n_{\text{run}}$ and repeated this entire process 20 times to report the mean and standard deviation of the averaged-SWD. As shown in Table LABEL:tab:var_swd, when $n_{\text{run}}=10$, the variance of the averaged-SWD is sufficiently small. We therefore adopt $n_{\text{run}}=10$ for all SWD computations in our experiments. In contrast, MMD is deterministic for a given kernel and does not exhibit this variance.
Metric Consistency Across Discretizations. A crucial property of a metric for functional data is its consistency with respect to the discretization of the function domain. The computed distance between two measures should be consistence regardless of the discretization for functions samples in two function datasets drawn from the two probability measure. That is, if we draw anthor new function samples $X_{\dagger}$ and $Y_{\dagger}$ from the same measures $\mu$ and $\nu$ but observe them on a different set of locations, we expect $\text{MMD}(X_{\dagger},Y_{\dagger})\approx\text{MMD}(X,Y)$ and $\text{SWD}(X_{\dagger},Y_{\dagger})\approx\text{SWD}(X,Y)$.
@@ C.2 Architectural Comparison to Other Neural Operators
The MINO architecture builds upon insights from previous work, including GINO and UPT (Li et al., 2023; Alkin et al., 2025), but introduces several key modifications to address their limitations in the context of functional generative tasks.
@@
Comparison with UPT. The UPT model uses a GNN for super-node pooling, a choice we identify as suboptimal for operator learning for three primary reasons. First, as the input discretization becomes finer, the GNN converges to a pointwise operator, which decouples it from the discretization convergence property required of a neural operator. Second, GNN can also be viewed as a low pass filter in the generative tasks, which has the same issue of GNO. Thrid, UPT relies on random subsampling to define super-nodes is ill-suited for adaptive meshes common in PDE tasks, where important regions with low grid density may be undersampled. For instance, on a dense mesh of the Earth, random subsampling would likely over-concentrate super-nodes in the polar regions that of less interest. Furthermore, UPT’s decoder uses a single-layer, perceiver-style cross-attention where the queries are derived only from the output positions. We found that for generative tasks, it is crucial to use a combination of the input function and its position as the query to the cross attention decoder and to use multiple attention layers as shown in the ablation study (Appendix E).
Comparison with Transolver. A key difference is that Transolver relies on a global receptive field, whereas our architecture combines both local and global information processing. Moreover, MINO more closely follows the modern, highly scalable, and efficient DiT architecture by replacing the grid-dependent “patchify” operation with our domain-agnostic GNO encoder. As demonstrated in our experiments, MINO is significantly more efficient than Transolver in both training and inference.
@@ Appendix E Ablation Study
Standard MINO-T $\mathbf{4.0\cdot 10^{-2}}$ $\mathbf{3.1\cdot 10^{-2}}$ $\mathbf{8.3\cdot 10^{-3}}$ $\mathbf{6.9\cdot 10^{-3}}$ $\mathbf{7.5\cdot 10^{-2}}$ $\mathbf{2.3\cdot 10^{-2}}$ $\mathbf{3.0\cdot 10^{-2}}$ $\mathbf{3.0\cdot 10^{-2}}$

### v3

@@ Acknowledgments
The authors would like to thank the TMLR reviewers and the action editor for their constructive feedback, and in particular reviewer fkny for the valuable suggestion regarding Table 8. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Science Foundations for Energy Earthshot under Award Number DE-SC0024705. ZER is supported by a fellowship from the David and Lucile Packard Foundation.
@@ A.1 Datasets description
Shallow Water. This dataset contains solutions to the shallow-water equations for a 2D radial dam-break scenario on a square domain, from PDEBench (Takamoto et al., 2022). Each of the 1,000 simulations has 1,000 time steps at $128\times 128$ resolution; we downsample spatially to $64\times 64$ for efficiency and treat each time step as an independent snapshot. We randomly select 30,000 snapshots for training and 5,000 for testing
Darcy Flow. This dataset contains steady-state solutions of 2D Darcy Flow over the unit square, obtained directly from the PDEBench benchmark (Takamoto et al., 2022). We downsample the original $128\times 128$ resolution to $64\times 64$. The dataset contains 10,000 samples and we split it into 8,000 samples for training and 2,000 for testing.
@@
Global Climate. We use the real-world global climate dataset from (Dupont et al., 2022), which contains global temperature measurements over the last 40 years. Each data sample is a function defined on a grid of $46\times 90$ evenly spaced latitudes and longitudes. Following the previous pipeline (Dupont et al., 2022), we convert the latitude-longitude pairs to Euclidean coordinates ($\mathbb{R}^{3}$) before passing them to the models. The dataset contains 9,676 training samples and 2,420 test samples.
@@ A.2 Baselines description
For all baseline models, we adopt their official implementations to ensure reproducibility; the corresponding code repositories are linked in the referenced papers. The key initialization hyperparameters for the baselines : GINO (Li et al., 2023), UPT (Alkin et al., 2024), Transolver (Wu et al., 2024), LNO (Wang & Wang, 2024), and FNO (Li et al., 2021)—are detailed in Table 6. Note that a slightly different configuration was used for the Global Climate experiment. For a complete understanding of all arguments, we refer the reader to the official repository for each respective model (github link provided in all referred papers)
@@
Latent Grid $32\times 32$ - - - -
@@ A.3 Hyperparameters for MINO
For the decoder, the single cross-attention block is followed by a multi-head self-attention (MHSA) block. Although this MHSA block is optional, we found that its inclusion slightly improves performance for generation task. To maintain the overall linear time complexity of MINO, the MHSA module can be replaced by any self-attention variant with linear complexity (like transolver layer).
MINO-T. The GNO maps input functions to a latent representation on a $16\times 16$ grid of query points, defined over the $0,1]^{2}$ domain. We set the GNO search radius to $0.07$, the latent dimension $L_{\text{dim}}$ to 256, and the number of attention heads to 4. The encoder consists of $M_{1}=5$ cross-attention blocks. For the Global Climate dataset, which is defined on a spherical manifold $\mathbb{S}^{2}$, we adjust the latent query positions to a $32\times 16$ spherical grid and increase the GNO radius to $0.2$ to account for the different coordinate system; other hyperparameters remain unchanged. The total parameter count for MINO-T is 21.5 M.
MINO-U. The GNO maps input functions to a latent representation on a $16\times 16$ grid of query points, defined over the $0,1]^{2}$ domain. We set the GNO search radius to $0.07$, the latent dimension $L_{\text{dim}}$ to 256, and the number of attention heads to 4. The encoder consists of $M_{1}=2$ cross-attention blocks. For its latent-space processor, we adopt a Diffusion U-Net architecture from the torchcfm library (Tong et al., 2024), which operates on the $[16,16]$ latent tensor with 64 channels, 1 residual block, and 4 attention heads for the processor. Same setting of MINO-T on the Global Climate dataset. The total parameter count for MINO-U is 19.2 M.
@@ A.4 Details for training and inference
Reference Measures. We choose $\mu_{0}$ as a Gaussian measure characterized by a Gaussian Process (GP) with a Matérn kernel. Unless otherwise specified, the function domain is $0,1]^{2}$ (or subset of it), and we use a kernel length scale of $0.01$ and a smoothness parameter of $0.5$. For the Global Climate dataset, the domain is a spherical manifold represented as a subset of $1,1]^{3}$, and the kernel length scale is adjusted to $0.05$. Furthermore, for the GP on the sphere, we explored two distance metrics for the kernel. The first uses the chordal distance (the Euclidean distance in the $\mathbb{R}^{3}$ embedding space), while the second employs the geodesic distance on $\mathbb{S}^{2}$, for which we leverage the GeometricKernels library (Mostowsky et al., 2024). Although both approaches yield a valid GP, the results presented in this paper are based on the chordal distance implementation. We selected this method for two primary reasons: its superior computational speed and to ensure consistency with other benchmarks. To facilitate further exploration, we will release the code repository with both implementations. The variance is fixed at 1 for all GPs.
Training Details. We train all models for 300 epochs using the AdamW optimizer with an initial learning rate of 1e-4. We employ a step learning rate scheduler that decays the learning rate by a gamma of $0.8$ every 25 epochs. The default batch size is 96. However, Transolver consumes significantly more GPU memory than other models. To accommodate it on a single NVIDIA RTX A6000 Ada GPU (48 GB memory), we reduced its batch size to 48 and, to maintain a comparable training iteration (duration), limited its training to 200 epochs. Despite these adjustments, Transolver still required approximately 1.76x more total GPU-hours than MINO-T and MINO-U, respectively. The only exception was the Global Climate experiment, where we explicitly matched the total GPU computation time of MINO variants to that of Transolver by extending their training epochs. All experiments reported in Table 4 and Table 3 were conducted three times, and we report the best performance (among the three) achieved for each model. To further strengthen our comparative analysis, Appendix G details an experiment with matched settings (parameter counts, batch size, epochs) that reveals MINO’s significant speedup and superior performance over Transolver.
Inference Details. To evaluate the models, we generate the same number of samples as contained in the test set for each dataset shown in Table 2. All samples are generated by solving the learned ODE numerically using the dopri5 solver from the torchdiffeq library (Chen et al., 2018), with an error tolerance set to 1e-5 for all experiments.
@@ Appendix B SWD and MMD as general metrics for functional generation tasks
[…]
@@
where $\theta$ is a random direction on the unit sphere $\mathbb{S}^{d-1}$. SWD converges to the exact p-Wasserstein distance for two probability measures as the number of projections goes to infinity. In our experiments, we use the official implementation from the POT library (Flamary et al., 2021). We use the Sliced Wasserstein-2 distance (p=2p=2(Bonneel et al., 2015), the time complexity for SWD is $\mathcal{O}(L\cdot N\cdot N_{in}+L\cdot N\log N)$, where $L$ is the number of projections.
@@
To validate this approach, we created two distinct Gaussian measures, characterized by GP, $\text{GP}_{1}$ (length scale=0.3, smoothness=1.5) and $\text{GP}_{2}$ (length scale=0.01, smoothness=0.5), on a $64\times 64$ regular grid. We drew 5,000 GP samples from each measure to create $X,Y$ and then calculated the averaged-SWD between them for varying $n_{\text{run}}$ and repeated this entire process 20 times to report the mean and standard deviation of the averaged-SWD. As shown in Table LABEL:tab:var_swd, when $n_{\text{run}}=10$, the variance of the averaged-SWD is sufficiently small. We therefore adopt $n_{\text{run}}=10$ for all SWD computations in our experiments. In contrast, MMD is deterministic for a given kernel and does not exhibit this variance.
Metric Consistency Across Discretizations. A crucial property of a metric for functional data is its consistency with respect to the discretization of the function domain. The computed distance between two measures should be consistent regardless of the discretization for functions samples in two function datasets drawn from the two probability measure. That is, if we draw new function samples $X_{\dagger}$ and $Y_{\dagger}$ from the same measures $\mu$ and $\nu$ but observe them on a different set of locations, we expect $\text{MMD}(X_{\dagger},Y_{\dagger})\approx\text{MMD}(X,Y)$ and $\text{SWD}(X_{\dagger},Y_{\dagger})\approx\text{SWD}(X,Y)$.
@@ C.2 Architectural Comparison to Other Neural Operators
The MINO architecture builds upon insights from previous work, including GINO and UPT (Li et al., 2023; Alkin et al., 2024), but introduces several key modifications to address their limitations in the context of functional generative tasks.
@@
Comparison with UPT. The UPT model uses a GNN for super-node pooling, a choice we identify as suboptimal for operator learning for three primary reasons. First, as the input discretization becomes finer, the GNN converges to a pointwise operator, which decouples it from the discretization convergence property required of a neural operator. Second, GNN can also be viewed as a low pass filter in the generative tasks, which has the same issue of GNO. Third, UPT relies on random subsampling to define super-nodes is ill-suited for adaptive meshes common in PDE tasks, where important regions with low grid density may be undersampled. For instance, on a dense mesh of the Earth, random subsampling would likely over-concentrate super-nodes in the polar regions that of less interest. Furthermore, UPT’s decoder uses a single-layer, perceiver-style (Jaegle et al., 2022) cross-attention where the queries are derived only from the output positions, which struggles to reconstruct the high-frequency information (Dong et al., 2021). We found that for generative tasks, it is crucial to use a combination of the input function and its position as the query to the cross attention decoder and to use multiple attention layers as shown in the ablation study (Appendix E). In Table 8, we present a detailed architectural comparison between MINO and GINO/UPT.
Table 8: Architectural Comparison: MINO vs. GINO / UPT
Feature MINO (Mesh-Informed Neural Operator) GINO / UPT Frameworks
Primary Goal Probabilistic Functional Generative Modeling (learning distributions via flow matching). Deterministic Physics Simulation: GINO—solving PDEs on varying geometries; UPT—scalable temporal simulation (latent rollouts).
Encoder Strategy Learned Resampling onto a Grid: a Graph Neural Operator (GNO) projects from $N_{\text{in}}$ irregular points to a fixed, regular $N_{\text{node}}$ grid. GINO: GNO maps points to a regular grid and augments them with geometric features (signed‐distance function). UPT: a GNN pools information from $N_{\text{in}}$ points onto a smaller, fixed set of $S$ “super-nodes.”
Latent Representation Structured Feature Map (Grid): a tensor of shape $(H\times W\times C)$, compatible to CNNs. GINO: structured feature map (grid). UPT: unstructured set of vectors (tokens) of shape $(N_{\text{tokens}}\times D_{\text{hidden}})$.
Core Processor SOTA diffusion-style U-Net (or Transformer). GINO: Fourier Neural Operator (FNO) for efficient global processing on the latent grid. UPT: Transformer (Approximator) for modeling temporal transitions between latent tokens.
Decoder Mechanism Cross-Attention with Input Bypass: Query (Q) comes from the original input function $f_{t}$ and its positions; Key/Value (KV) come from the processed latent state. Prevents an information bottleneck and preserves high-frequency details. No Input Bypass (Information Bottleneck): GINO projects directly from latent grid to query points. UPT uses Perceiver-style cross-attention where Q is derived from output positions only. All information must be reconstructed solely from the compressed latent state.
Uncertainty Inherent & Computationally Cheap: being generative, the model learns a distribution; Monte-Carlo dropout can provide epistemic uncertainty with a single trained network. Computationally Expensive: uncertainty is not a primary target; estimating it usually requires training an ensemble of models (multiple runs with different seeds).
Comparison with Transolver. Key differences lie in the receptive field and latent‑space design of MINO. Transolver employs a purely global receptive field, whereas our architecture integrates both local and global information processing. In addition, MINO follows the modern, highly scalable, and efficient DiT paradigm by replacing the grid‑dependent “patchify” operation with a domain‑agnostic GNO encoder. Finally, because the slicing and de‑slicing operations in Transolver act on every point, and these operations happen in every Transolver layer, which can be inefficient. In contrast, MINO performs encoding and decoding only once and can be viewed as a reduced‑order operator; its decoder is shallow, consisting of one cross‑attention block followed by a single self‑attention block. As shown in our experiments, MINO is markedly more efficient than Transolver in both training and inference.
@@ Appendix E Ablation Study
Standard MINO-T $\mathbf{4.0\cdot 10^{-2}}$ $\mathbf{3.6\cdot 10^{-2}}$ $\mathbf{9.8\cdot 10^{-3}}$ $\mathbf{8.7\cdot 10^{-3}}$ $\mathbf{8.9\cdot 10^{-2}}$ $\mathbf{3.4\cdot 10^{-2}}$ $\mathbf{2.9\cdot 10^{-2}}$ $\mathbf{2.6\cdot 10^{-2}}$
@@
Appendix F Empirical Analysis of Scaling and Stability
To evaluate performance consistency, we train both MINO-U and MINO-T with eight different random seeds on the Navier-Stokes and Cylinder-Flow datasets. The results, summarized in Table 10, show that MINO-U consistently outperforms MINO-T across both datasets and metrics. We also note that performance variance across seeds is slightly higher on the Navier-Stokes task compared to the Cylinder-Flow task.
We next study data-scaling on Navier–Stokes benchmark by training on progressively larger subsets of the full 30 k-sample dataset and evaluating on the same test split. Table LABEL:tab:scaling shows steady performance gains as the training set grows, with improvements tapering off once roughly 18 k samples are reached.
Table 10: Performance comparison over eight random seeds. We report the mean ± standard deviation (in the parentheses) for each metric. Lower is better.
Dataset $\rightarrow$ Navier Stokes Cylinder Flow
Model $\downarrow$ Metric $\rightarrow$ SWD MMD SWD MMD
MINO-T $3.7\cdot 10^{-2}~(6.8\cdot 10^{-3})$ $3.1\cdot 10^{-2}~(6.0\cdot 10^{-3})$ $2.9\cdot 10^{-2}~(2.4\cdot 10^{-3})$ $2.6\cdot 10^{-2}~(3.0\cdot 10^{-3})$
MINO-U $3.3\cdot 10^{-2}~(6.3\cdot 10^{-3})$ $2.7\cdot 10^{-2}~(7.5\cdot 10^{-3})$ $2.1\cdot 10^{-2}~(2.0\cdot 10^{-3})$ $1.9\cdot 10^{-2}~(2.8\cdot 10^{-3})$
Table 11: Performance of MINO-U when trained on a subset of the Navier-Stokes dataset.
Metrics $\downarrow$ Samples $\rightarrow$ 6,000 12,000 18,000 24,000 30,000
SWD $4.3\cdot 10^{-2}$ $3.9\cdot 10^{-2}$ $2.8\cdot 10^{-2}$ $3.0\cdot 10^{-2}$ $2.8\cdot 10^{-2}$
MMD $4.0\cdot 10^{-2}$ $3.2\cdot 10^{-2}$ $2.2\cdot 10^{-2}$ $3.1\cdot 10^{-2}$ $1.9\cdot 10^{-2}$
Appendix G Comparison with Transolver under Identical Settings
To ensure a fair comparison with Transolver, we reduce the latent dimension of MINO from 256 to 192, and use the same training configuration (200 epochs, batch size 48). This results in smaller models—MINO-U (S) (13.5M parameters) and MINO-T (S) (12.1M)—compared to Transolver (15.0M). Notably, our models are significantly faster, achieving a 3.1x speedup in training and a 3.5x speedup in inference while maintaining superior performances.
Table 12: Performance comparison with Transolver under identical settings. Best Performance in bold
Dataset $\rightarrow$ Navier Stokes Mesh GP Cylinder Flow
Model $\downarrow$ Metric $\rightarrow$ SWD MMD SWD MMD SWD MMD
Transolver $5.3\cdot 10^{-2}$ $5.1\cdot 10^{-2}$ $9.1\cdot 10^{-2}$ $4.6\cdot 10^{-2}$ $2.5\cdot 10^{-2}$ $2.4\cdot 10^{-2}$
MINO-T (S) $4.4\cdot 10^{-2}$ $4.0\cdot 10^{-2}$ $7.9\cdot 10^{-2}$ $\mathbf{2.8\cdot 10^{-2}}$ $2.7\cdot 10^{-2}$ $2.6\cdot 10^{-2}$
MINO-U (S) $\mathbf{2.9\cdot 10^{-2}}$ $\mathbf{2.5\cdot 10^{-2}}$ $\mathbf{6.1\cdot 10^{-2}}$ $3.0\cdot 10^{-2}$ $\mathbf{2.5\cdot 10^{-2}}$ $\mathbf{2.3\cdot 10^{-2}}$